P27. Given the following data for TMC during January 200X

Cost of goods sold

120,000

Material purchased

40,000

Finished goods were decreased by

3,000

Work in process inventory increased by

4,000

Direct material inventory decreased by

2,000

Factory overhead applied based on

3

times of direct labor costs

Find 1) Conversion costs 2) Prime costs
Answer #1

Direct material used = 40000+2000 = 42000

Cost of goods manufactured = 120000-3000 = 117000

Total manufacturing cost = 117000+4000 = 121000

1) Conversion Cost = 121000-42000 = 79000

2) Direct labor =79000/4 = 19750

Prime cost = 19750+42000 = 61750